OpenStudy (anonymous):

Mac set up a refreshment stand at the dog show to raise money for his class trip. He is selling doughnuts for $2 and bagels with cream cheese for $3. He used the equation m = 2d + 3b, where m represents the money raised, d represents the number of doughnuts, and b represents the number of bagels with cream cheese. If he sells 80 doughnuts, how many bagels must he sell to raise $310 for his class trip?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

310=2(80)+3B 310=160+3B 310-160=3B 150=3B 150/3=B 50=B